40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 12002     Entry time: Mon Feb 22 13:56:52 2016     In reply to: 11999     Reply to this: 12006
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap -reflected beam from ETM aligned 

I tried aligning the green beam, elliptical as it is, to the arm by using the various steering mirrors after the doubling oven. The following was done:

  1. Eric and I aligned the beam through the green Faraday - we levelled the beam using an iris to check the beam height immediately after the Faraday and a little further along the beam propagation direction.
  2. We checked that the beam is reasonably centered on all the lenses. We changed the lens holder for one of the lenses from a Thorlabs model to a Newport model, so as to get the lens to the correct height such that the green beam was roughly centered on it. 
  3. I then tweaked the alignment of the steering mirrors until the reflected beam from the ETM roughly coincided with the input beam. The return beam is getting clipped slightly on the way back through the green Faraday, so some more alignment needs to be done. However, given the ITMX situation, I can't align the arm to IR, so I'm holding off on further alignment for now...
ELOG V3.1.3-